(1/3)^2 + 1/4 - (4/3)^2 * (1/2)^3
Answer by stanbon(75887)

(1/3)^2 + 1/4 - (4/3)^2 * (1/2)^3
-----
= (1/9) + (1/4) - (16/9)*(1/8)
= (13/36) - (8/36)
= 5/36
=============
Cheers,
Stan H.